# Xiangqi Wang1, Xiuying Chen2, Mohamed Elhoseiny3, Xiangliang Zhang1,†

### v1

@@ Appendix E Error Study
Appendix F Prompt Template
Prompt Template: Generate Aspects [System]
You are an AI assistant specializing in designing prompts to test Large Vision-Language Models (LVLMs). Your task is to create meticulously {aspect_count} fined-grained aspects that evaluate LVLMs basic understanding of images.
[Background] Large Vision-Language Models are AI systems capable of understanding and analyzing images. Testing these models across various competencies is crucial for assessing their performance, limitations, and potential biases. The aspects you create will be used to challenge and evaluate LVMs.
[Instruction]
1.Basic Understanding: This involves recognizing and identifying individual objects, characters, and scenes within an image. It includes tasks like detecting the presence of specific items (e.g., cars, trees, people), distinguishing between different types of objects, and understanding the general context of the scene (e.g., a park, a city street). The goal is to accurately label all relevant elements in the image, providing a foundation for more advanced analysis.
2.Come up with 4 general aspects according to the basic understanding.
3.Then Create 6 fined-grained aspects within the basic understanding for each general aspect, do not go beyond. You can consider the definition of the basic understanding above. 4. Please list the aspects without using numbered lists. 5. Let’s think step by step.
[Output Format]
Please strictly respond in the following format:
General Aspect: [Aspect]
Fined-grained Aspect: [Aspect]
Introduction: [Introduction]
Figure 22: Prompt template for generating aspects. Here we use the task of basic understanding as an example.
Prompt Template: Generate Guideline [System]
You are an advanced AI simulation assistant specializing in crafting precise prompts for image generation models.
[Instruction] I will provide you with an aspect for image generation. Your task is to create a detailed instruction on how to incorporate this aspect into an image generation prompt. This instruction should guide an AI to produce a prompt that will result in a image.
Here are two examples to guide you:
…If the aspect is Foreground vs. Background, then the prompt should have what’s in the foreground and what’s in the background.
…If the aspect is Relative Size Estimation, then you need to include the size of one thing compared to another.
The aspect we’ll focus on is: {aspect}, and the introduction is: {introduction}
When crafting your instruction, consider the following:
1. Be specific about how the aspect should be represented visually, similar to the examples provided.
2. Provide clear guidelines on how to balance different elements within the image.
3. Include tips on avoiding common pitfalls or misinterpretations related to this aspect.
Remember, the goal is to instruct an AI on creating a prompt that will generate a single, coherent image. Your instruction should be comprehensive enough to ensure the final prompt will produce a high-quality, well-integrated result.
[Output Format]
Aspect: {aspect}
Introduction: {introduction}
Guidance: [Provide your instruction directly, written in a clear, authoritative tone. Do not include any explanations, disclaimers, or additional commentary outside of the instruction itself. Similar to the example given above, do not divide it into sections.]
Figure 23: Prompt template for generating guideline.
Prompt Template: Generate Image Description [System]
You are an AI assistant tasked with converting user inputs and their descriptions into suitable prompts for a text-to-image model. These prompts will generate images to test the capabilities of large vision language models (LVLMs).
[Background] Large Vision Language Models (LVLMs) are AI systems proficient in interpreting and analyzing images. Evaluating these models across different competencies is essential to understanding their performance, limitations, and potential biases. The prompts you create will be used to generate images through text-to-image models, which will then be used to challenge and evaluate LVLMs.
[Instruction]
1.Carefully analyze the given aspect: {aspect}, its introduction: {introduction} and prompt generation guidance: {guidance}
2.Generate a suitable prompt based on the provided aspect and introduction for the text-to-image model to create an image. Ensure that the prompt is composed of simple phrases, avoiding overly complex descriptions, and is clear enough. If you deem the description irrelevant to the test content, do not generate a related prompt.
3.Consider including elements that might be particularly challenging for LVMs, such as unusual combinations, abstract concepts, or subtle details.
4.We categorize the difficulty of prompts into easy, medium, and hard: ...
5.Provide one overarching topic word that encapsulates the essence of your description.
6.List 4-6 key words that are closely related to your description and crucial for understanding the image.
7.Avoid using the following words in your new description: {used_words_str}
8.The required difficulty level is: {level}
9.Please use clear and accurate words, clear logic flow, do not use too abstract words. Word Choice: Word choice matters. More specific synonyms work better in many circumstances. Instead of big, try tiny, huge, gigantic, enormous, or immense. Plural words and Collective Nouns: Plural words leave a lot to chance. Try specific numbers. "Three cats" is more specific than "cats." Collective nouns also work, “flock of birds” instead of "birds.” Focus on What You Want: It is better to describe what you want instead of what you don’t want. If you ask for a party with “no cake,” your image will probably include a cake. Try to be clear about any context or details that are important to you. Think about: Subject: person, animal, character, location, object Medium: photo, painting, illustration, sculpture, doodle, tapestry Environment: indoors, outdoors, on the moon, underwater, in the city Lighting: soft, ambient, overcast, neon, studio lights Color: vibrant, muted, bright, monochromatic, colorful, black and white, pastel Mood: sedate, calm, raucous, energetic Composition: portrait, headshot, closeup, birds-eye view But don’t write it directly in colon form, but express it normally in a sentence. ]
[Output Format]
Please strictly respond in the following format:
Aspect: {aspect}
Prompt: [Your detailed image description]
Topic word: [One word that captures the essence of the description]
Key word: [Word1, Word2, Word3,...]
Figure 24: Prompt template for generating image descriptions with difficulty grading.
Prompt Template: Generate Q&A
[System] You are an AI assistant tasked with converting user inputs and their descriptions into suitable questions to test the Large Vision-Language Model’s (LVLM) abilities in given aspects.
[Background] Large Vision-Language Models (LVLMs) are AI systems proficient in interpreting and analyzing images. Evaluating these models across different competencies is essential to understanding their performance, limitations, and potential biases. We will provide you with a prompt to generate an image, which will create a specific image. You can then formulate questions about this image based on the prompt. The questions you create will be used to challenge and evaluate LVLMs based on generated images.
[Instruction]
1.Carefully analyze the given aspect and its Introduction: Aspect:{aspect}.
2.Generate a suitable question based on the provided image generation prompt, and aspect to test the LVLM’s ability in the given aspect. Ensure that the question is related to the prompt of the image and is of moderate difficulty.
3.We categorize the difficulty of questions into easy, medium, and hard. For easy difficulty, please formulate questions based solely on very simple details from the image generation prompt, ensuring they adhere to the scope of the aspect in question. For medium difficulty, ensure the question is challenging but not overly complex, involving common scenes and requiring some level of inference or detailed observation. For hard difficulty, consider incorporating elements that may be particularly challenging for LVLMs, such as unusual combinations or subtle details, while keeping the question clear and relevant, and ensure it is more demanding than the medium level.
4.Avoid using overly complicated language or details unrelated to the image in the questions.
5. When generating problems of different difficulty, please combine the current specific aspect.
6.Due to potential discrepancies in image generation, we have detected the following errors:{elements}. Please avoid referencing these elements in your questions. If the prompt for generating the image does not describe in detail what the specific looks like, please do not ask related questions.
7.The required difficulty level is:{level}
8.Please generate a multiple-choice question, which can either be a four-option single-choice question or a true/false question. If it is a true/false question, the options should be A. True B. False.
Image generation prompt:{prompt}
9. The answers in the options need to be differentiated to a certain extent. There cannot be a situation where multiple options meet the requirements of the question. There can only be one answer that meets the question.
Aspect:{aspect}
[Output Format]
Please directly output the generated question in the following JSON format:
"question": "[your question]",
"options": {
"A": "[Option A]",
"B": "[Option B]",
"C": "[Option C]",
"D": "[Option D]"
"reference_answer": "A or B or C or D"
Without any other information.
Figure 25: Prompt template for generating questions.
Prompt Template: Answer [System]
In order to test your ability with pictures, we have a question about aspect area. Please answerbased on your knowledge in this area and your understanding of pictures.
Given the image below, answer the questions: question based on the image.
Please give the final answer strictly follow the format [[A]] (Srtictly add [[ ]] to the choice, and the content in the brackets should be the choice such as A, B, C, D) and provide a brief explanation of your answer. Directly output your answer in this format and give a brief explanation.
Figure 26: Prompt template for LVLM to answer.

### v3

@@ Appendix E Notations
Appendix F Error Study
Through extensive experimental analysis, we have categorized the common problems encountered by LVLMs in VQA tasks into two main types: image comprehension errors and image reasoning errors. Regarding the first category, LVLMs often fail to truly understand the details in an image. For instance, in Figure 15, the model failed to accurately identify the performer’s attire. In Figure 17, from the model’s explanation, it is evident that the focus of its understanding was placed on the text rather than the image. Option B mentions a swimming pool, option C references a helipad, and option D refers to a health center—none of which are present in the image. In Figure 19, the model failed to correctly identify the magician’s facial expression as a smile. Therefore, it cannot be associated with calmness or scheming; instead, it represents a confident smile. These errors demonstrate the model’s inadequate comprehension of image details.
Image reasoning errors occur when models accurately perceive the image content but falter in their reasoning process, leading to incorrect answers. For instance, in Figure 16, the model correctly identified the blue whale and the small boat in the image. However, based on its explanation, it relied more on its prior knowledge rather than the image itself, leading to an incorrect answer. In Figure 18, model accurately recognized the excitement and joy on the little girl’s face but misinterpreted the intent. From the image, it is clear that the woman is holding a complete ice cream cone and is about to hand it to the girl, indicating that the girl desires the ice cream.
Figure 15: Error study of Llama-3.2-90B-Vision under basic understanding.
Figure 16: Error study of Claude-3-Haiku under spatial understanding.
Figure 17: Error study of Gemini-1.5-Flash under semantic understanding.
Figure 18: Error study of Qwen2-VL under reasoning capacity.
Figure 19: Error study of GLM-4v-Plus under atmospheric understanding.
Appendix G Case Study
Figure 20: Case Study 1.
Figure 21: Case Study 2.
Figure 22: Case Study 3.
Figure 23: Case Study 4.
Figure 24: Case Study 5.
Figure 25: Case Study 6.
Figure 26: Case Study 7.
Figure 27: Case Study 8.
Figure 28: Case Study 9.
Figure 29: Case Study 10.
Figure 30: Case Study 11.
Figure 31: Case Study 12.
Figure 32: Case Study 13.
Appendix H Prompt
Aspect Generation You are an AI assistant specializing in designing prompts to test Large Vision-Language Models (LVLMs). Your task is to create meticulously aspect_count fined-grained aspects that evaluate LVLMs basic understanding of images. Large Vision-Language Models are AI systems capable of understanding and analyzing images. Testing these models across various competencies is crucial for assessing their performance, limitations, and potential biases. The aspects you create will be used to challenge and evaluate LVMs. 1. Basic Understanding: This involves recognizing and identifying individual objects, characters, and scenes within an image. It includes tasks like detecting the presence of specific items (e.g., cars, trees, people), distinguishing between different types of objects, and understanding the general context of the scene (e.g., a park, a city street). The goal is to accurately label all relevant elements in the image, providing a foundation for more advanced analysis. 2. The aspects you generate must test the understanding of a single image, not multiple images, e.g. multiple perspectives. 3. Come up with 4 general aspects according to the basic understanding. 4. Then Create 6 fined-grained aspects within the basic understanding for each general aspect, do not go beyond. You can consider the definition of the basic understanding above. 5. List the aspects without using numbered lists. 6. Let’s think step by step. Please strictly respond in the following format: General Aspect: [Aspect] Fined-grained Aspect: [Aspect] Introduction: [Introduction]
Image Description Generation-1 You are an AI assistant tasked with converting user inputs and their descriptions into suitable prompts for a text-to-image model. These prompts will generate images to test the capabilities of large vision language models (LVLMs). Large Vision Language Models (LVLMs) are AI systems proficient in interpreting and analyzing images. Evaluating these models across different competencies is essential to understanding their performance, limitations, and potential biases. The prompts you create will be used to generate images through text-to-image models, which will then be used to challenge and evaluate LVLMs. 1. Carefully follow the given aspect: aspect, its introduction: introduction. 2. Generate a suitable prompt based on the provided aspect and introduction for the diffusion model to create an image. Ensure that the prompt is composed of simple phrases, avoiding overly complex descriptions, and is clear enough. If you deem the description irrelevant to the test content, do not generate a related prompt. 3. Consider including elements that might be particularly challenging for LVMs, such as unusual combinations, abstract concepts, or subtle details. 4. Details about different difficulty levels: - Easy Difficulty - Characteristics: Focus on a single subject but with minor additional complexity to test precision in details or context. - Purpose: Establish baseline capability with subtle challenges, such as fine texture, simple interactions, or slight variations in lighting. - Description Requirements: A single object or entity with some basic contextual details or additional features, placed against a minimally distracting background. - Examples:- “A perfectly polished red apple with a small leaf attached, placed on a slightly reflective white surface.” - “A blue balloon with a slightly wrinkled texture, tied to a thin white string, floating against a pale sky with soft clouds.” - Key Points: Test finer details (e.g., texture, reflection) while still keeping the composition simple and clear. - Medium Difficulty - Characteristics: Scenes involve multiple elements or interactive settings that require nuanced spatial arrangement and accurate relationships between objects. - Purpose: Evaluate the ability to generate cohesive, moderately dynamic scenes with layered realism and a stronger sense of depth. - Description Requirements: Include multiple objects interacting naturally in a believable environment, with more intricate details and subtle light or shadow effects. - Examples: - “A steaming cup of coffee on a wooden table, with a half-eaten croissant beside it, morning sunlight casting soft shadows through a lace curtain in the background.” - “A golden retriever running on a sandy beach, splashing water as it chases a bright orange ball, with distant waves and a partly cloudy sky.” - Key Points: Incorporate realistic environmental elements and ensure spatial coherence, emphasizing interactions and secondary details (e.g., shadows, water splashes). - Hard Difficulty - Characteristics: Scenes incorporate high complexity with multiple interdependent elements, challenging perspectives, or dynamic and intricate lighting or textural effects. - Purpose: Push the limits of rendering capability to handle advanced relationships, environmental effects, and challenging compositions. - Description Requirements: The scene must include 3-4 main elements or a combination of dynamic features, such as motion, light interplay, or atmospheric conditions, while maintaining clarity and logic. - Examples: - “A raindrop-streaked window reflecting the interior of a cozy room, with a black cat sitting on the windowsill and a glowing city skyline visible through the glass, all under the warm hues of a sunset.” - “A knight in shining armor standing on a cliff edge, overlooking a stormy sea, with bolts of lightning illuminating the dark clouds and waves crashing against jagged rocks below.” - Key Points: Highlight challenges such as complex reflections, dynamic light, or multi-element interactions, ensuring visual harmony and detailed textures.
Image Description Generation-2 5. Provide one overarching topic word that encapsulates the essence of your description. 6. List 4-6 key words that are closely related to your description and crucial for understanding the image. 7. Avoid using the following words in your new description: used_words_str 8. The required difficulty level is: level 9. Please use clear and accurate words, clear logic flow, do not use too abstract words. The length of the generated sentences needs to be consistent with the examples. 10. Just output the image description used to generate the image, don’t mention anything else Please strictly respond in the following format: Aspect: aspect Prompt: [Your detailed image description] Topic word: [One word that captures the essence of the description] Key word: [Word1, Word2, Word3,…]
Align Question Generation Given the image descriptions:description, generate six questions (True/False or Multiple choice) with only one correct choice that verifies if the image description is correct. Classify each concept into a type (object, human, animal, food, activity, attribute, counting, color, material, spatial, location, shape, other), and then generate a question for each type. Here’s some examples: ”’Description: A man posing for a selfie in a jacket and bow tie. Entities: man, selfie, jacket, bow tie Activities: posing Colors: Counting: Other attributes: Questions and answers are below: About man (human): Q: is this a man? Choices: yes, no A: yes Q: who is posing for a selfie? Choices: man, woman, boy, girl A: man About selfie (activity): Q: is the man taking a selfie? Choices: yes, no A: yes Q: what type of photo is the person taking? Choices: selfie, landscape, sports, portrait A: selfie About jacket (object): Q: is the man wearing a jacket? Choices: yes, no A: yes Q: what is the man wearing? Choices:jacket, t-shirt, tuxedo, swearter A: jacket About bow tie (object): Q: is the man wearing a bow tie? Choices: yes, no A: yes Q: is the man wearing a bow tie or a neck tie? Choices: bow tie, neck tie, cravat, bolo tie A: bow tie About posing (activity): Q: is the man posing for the selfie? Choices: yes, no A: yes Q: what is the man doing besides taking the selfie? Choices: posing, waving, nothing, shaking A: posing Description: A horse and several cows feed on hay. Entities: horse, cows, hay Activities: feed on Colors: Counting: several Other attributes: Questions and answers are below: About horse (animal): Q: is there a horse? Choices: yes, no A: yes About cows (animal): Q: are there cows? Choices: yes, no A: yes About hay (object): Q: is there hay? Choices: yes, no A: yes Q: what is the horse and cows feeding on? Choices: hay, grass, leaves, twigs A: hay About feed on (activity): Q: are the horse and cows feeding on hay? Choices: yes, no A: yes About several (counting): Q: are there several cows? Choices: yes, no A: yes ”’ And finally respond in the following format: {{ "caption": "description", "question": "Your question here", "choices": [yes or no], "answer": "give your correct answer", "element_type": "Type of the element", "element": "Element name" }}, {{ "caption": "description", "question": "Your question here", "choices": [yes or no], "answer": "give your correct answer", "element_type": "Type of the element", "element": "Element name" }},
Align Answer Generation Given the image below, answer the questions: question from the choice: choices based on the image. And directly give the answer. {{ "answer":"yes or no" }}
Question Generation-1 You are an AI assistant tasked with converting user inputs and their descriptions into suitable questions to test the Large Vision Model’s (LVM) abilities in given aspects. Large Vision Models (LVMs) are AI systems proficient in interpreting and analyzing images. Evaluating these models across different competencies is essential to understanding their performance, limitations, and potential biases. We will provide you with a prompt to generate an image, which will create a specific image. You can then formulate questions about this image based on the prompt. The questions you create will be used to challenge and evaluate LVMs based on generated images. 1. Carefully analyze the given aspect and its Introduction: Aspect:aspect. 2. Generate a suitable question based on the provided image description to test the LVM’s ability in the given aspect. 3. We categorize the difficulty of questions into easy, medium, and hard: - Easy Difficulty:Focus on questions that require the identification of simple, prominent, and explicit details within the image. These questions should be straightforward, relying solely on basic observation without the need for inference or interpretation. For example, you might ask about the color of a specific object, the presence of a single item, or the shape of an easily recognizable feature. The key is to keep the questions direct and simple, ensuring that the answer is obvious and immediately visible in the image. - Medium Difficulty:Design questions that necessitate a moderate level of observation and inference. These questions should involve understanding relationships between elements, recognizing interactions, or identifying less prominent features that are still clear but not immediately obvious. Examples could include questions about the relative position of objects, identifying an action taking place, or understanding the context of a scene. The goal is to require some level of thought beyond basic observation, challenging the model to understand the scene’s composition or narrative without being overly complex. - Hard Difficulty:Create questions that require the model to notice and interpret more detailed aspects of the image. These questions should involve recognizing multiple elements working together, understanding more complex interactions, or identifying details that are present but not immediately obvious. For example, you might ask about the positioning of objects relative to each other in a more crowded scene, subtle changes in lighting or color that affect the appearance of objects, or identifying an element that is not the main focus but still visible in the background. The aim is to challenge the model to go beyond surface-level details, but without making the task too abstract or overly difficult.
Question Generation-2 4. Avoid using overly complicated language or details unrelated to the image in the questions. 5. When generating problems of different difficulty, please combine the current specific aspect. 6. Due to potential discrepancies in image generation, we have detected the following errors: elements. Please avoid referencing these elements in your questions. If the prompt for generating the image does not describe in detail what the specific looks like, please do not ask related questions. For example, if the prompt mentions a forest with glowing plants but does not specify how many there are, please do not ask a question about counting the number of glowing plants. 7. The required difficulty level is: level 8. Please generate a multiple-choice question, which is four-option single-choice question. 9. The answers in the options need to be differentiated to a certain extent. There cannot be a situation where multiple options meet the requirements of the question. There can only be one answer that meets the question. Image generation prompt: prompt Aspect: aspect Please directly output the generated question in the following JSON format: "question": "[your question]", "options": "A": "[Option A]", "B": "[Option B]", "C": "[Option C]", "D": "[Option D]" "reference_answer": "A or B or C or D" Without any other information and remember only one option in the reference answer.
Option Adjustment You are an ai assistant tasked with answering questions based on the given picture description without given the image. - Answer the questions based on your knowledge. - Please note that some incorrect answers are provided below. You must not make the same mistakes - Your answer needs to be semantically distinct from the given incorrect answer. - Don’t say you can’t see the image, just answer based on your knowledge. - Don’t generate overly lengthy answers, keep them concise and to the point. - The answer you generate needs to be factually different from the given incorrect answer. - Try to use straightforward words instead of being too abstract or vague. question:question wrong answers:wrong_answer Directly give you answer, don’t add thinkings or other information.
Answer Generation In order to test your ability with pictures, we have a question about aspect area. Please answer based on your knowledge in this area and your understanding of pictures. Given the image below, answer the questions: question based on the image. Please give the final answer strictly follow the format [[A]] (Strictly add [[ ]] to the choice, and the content in the brackets should be the choice such as A, B, C, D) and provide a brief explanation of your answer. Directly output your answer in this format and give a brief explanation. If you cannot read the picture, just answer based on your text ability.